### v3

@@ 5.2 Our Formal Algorithm
We now describe our formal algorithm. From the output of Algorithm 1, we characterize one optimal solution. If the sample size $n$ is used in Algorithm 1, we denote by $\mathcal{I}_{n}$ and $\mathcal{J}_{n}$ the output of Algorithm 1. To be specific, we have $\bm{q}^{*}_{\mathcal{I}_{n}^{c}}=0$ and the non-zero elements $\bm{q}^{*}_{\mathcal{I}_{n}}$ can be given as the solution to
@@
However, in practice, both the matrices $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ are unknown. We aim to use the samples to learn the matrices $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ such that the $\bm{q}^{*}_{\mathcal{I}_{n}}$ can also be determined. Our formal algorithm is given in Algorithm 2. The steps 3-8 in Algorithm 2 is to use Algorithm 1 as a subroutine to identify the set $\mathcal{I}$ and $\mathcal{J}$ that satisfy the conditions in Lemma 3.1. We can show that as long as $n\geq N_{0}^{\prime}$, where $N_{0}^{\prime}$ is a threshold that depends on the problem parameters, Algorithm 1 correctly obtains the set $\mathcal{I}$ and $\mathcal{J}$ satisfying the conditions in Lemma 3.1. Therefore, we exponentially increase the value of $n$ as input to Algorithm 1 to reach $N_{0}^{\prime}$.
@@
1:Input: the number of samples $N$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$.
2:Initialize $\mathcal{F}_{1}=\emptyset$, $\bm{\alpha}^{1}=N\cdot\bm{\alpha}$ and $\bm{\mu}^{1}=N\cdot\bm{\mu}$.
3:for $n=1,\dots,N$ do
4:     if $n=2^{m}$ for an integer $m$ then
5:         Obtain the output $\mathcal{I}_{n}$ and $\mathcal{J}_{n}$ from Algorithm 1 with the input $\mathcal{F}_{n}$.
6:     else
7:         set $\mathcal{I}_{n}=\mathcal{I}_{n-1}$ and $\mathcal{J}_{n}=\mathcal{J}_{n-1}$.
8:     end if
9:     Construct estimates $\bar{C}^{n}(\mathcal{J}_{n},\mathcal{I}_{n})$ and $\bar{B}^{n}(:,\mathcal{I}_{n})$ using the sample set $\mathcal{F}_{n}$.
10:     Construct a solution $\tilde{\bm{q}}^{n}$ such that $\tilde{\bm{q}}^{n}_{\mathcal{I}_{n}^{c}}=0$ and $\tilde{\bm{q}}^{*}_{\mathcal{I}_{n}}$ is the solution to
@@
11:     Project $\tilde{\bm{q}}^{n}$ to the set $\{\bm{q}:\|\bm{q}\|_{1}\leq 2\}$ to obtain $\bm{q}^{n}$.
12:     For each $(s,a)\in\mathcal{I}_{n}\subset\mathcal{S}\times\mathcal{A}$, we query the model $\mathcal{M}$ to obtain a sample of the reward $r^{n}(s,a)$ and the costs $c^{n}_{k}(s,a)$ for each $k\in\mathcal{J}_{n}\subset[K]$, as well as the state transition $s^{n}(s,a)$.
13:     Update $\mathcal{F}_{n+1}=\mathcal{F}_{n}\cup\left\{r^{n}(s,a),c^{n}_{k}(s,a),s^{n}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A},\forall k\in[K]\right\}$.
14:     Denote by $\bm{c}^{n}(s,a)=(c^{n}_{k}(s,a))_{\forall k\in\mathcal{J}_{n}}$ and do the update:
$\bm{\alpha}^{n+1}_{\mathcal{J}_{n}}=\bm{\alpha}^{n}_{\mathcal{J}_{n}}-\sum_{(s,a)\in\mathcal{I}_{n}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a).$ (24)
15:     Do the update:
$\bm{\mu}^{n+1}(s^{\prime})=\bm{\mu}^{n}(s^{\prime})-\sum_{(s,a)\in\mathcal{I}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall s^{\prime}\in\mathcal{S}.$ (25)
@@
16:end for
17:Output: $\bm{q}^{1},\dots,\bm{q}^{N}$.
A crucial element in Algorithm 2 (step 10) is that we adaptively update the value of $\bm{\alpha}^{n}_{\mathcal{J}_{n-1}}$ and $\bm{\mu}^{n}(s^{\prime})$ as in (24) and (25). We then use the updated $\bm{\alpha}^{n}_{\mathcal{J}_{n-1}}$ and $\bm{\mu}^{n}(s^{\prime})$ to obtain the value of $\tilde{\bm{q}}^{n}_{\mathcal{I}_{n}}$ as in (23). Such an algorithmic design has the following natural interpretation and is the key for us to achieve a logarithmic-level regret. Note that if we adopt a static way to compute $\bm{q}^{n}$, for example, we simply solve (22) with the matrices $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ replaced by their estimates, then the gap between $\bm{q}^{n}$ and $\bm{q}^{*}$ will be at the same order of the estimation error of $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$, which is $O(1/\sqrt{n})$. Thus, the gap between $\frac{\sum_{n=1}^{N}\bm{q}^{n}}{N}$ and $\bm{q}^{*}$ will be at the order of $O(1/\sqrt{N})$, which translates into $O(\sqrt{N})$ regret bound. Instead, we use $\bm{\alpha}^{n}_{\mathcal{J}_{n-1}}$ and $\bm{\mu}^{n}(s^{\prime})$ that are adaptive to the historical information. To be specific, if for one $n\in[N]$ and one binding constraint $k\in\mathcal{J}$, the constraint value under the action $\sum_{n^{\prime}=1}^{n}\bm{q}^{n^{\prime}}/n$ is below the target $\alpha_{k}$, we know that $\alpha^{n}_{k}/(N-n+1)$ is greater than $\alpha_{k}$, which results in a greater constraint value for $\tilde{\bm{q}}^{n+1}$ (it is in fact $\alpha^{n}_{k}/(N-n+1)$ beyond the target $\alpha_{k}$). In this way, any gap between the real-time constraint value and the target will be self-corrected in the next period as we adaptively obtain $\tilde{\bm{q}}^{n+1}$. Therefore, such an adaptive design can re-adjust the possible constraint violation by itself, which results in a regret bound of a lower order. In step 11, we further project $\tilde{\bm{q}}^{n+1}$ to the set $\{\bm{q}\geq 0:\|\bm{q}\|_{1}\leq 2\}$ to obtain $\bm{q}^{n+1}$. This prevents $\tilde{\bm{q}}^{n}$ from behaving ill when $n$ is not large and the estimates of $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ are not accurate enough. We can show that when $n$ is large enough, $\tilde{\bm{q}}^{n+1}$ automatically belongs to the set $\{\bm{q}\geq 0:\|\bm{q}\|_{1}\leq 2\}$.
Remark. Note that the adaptive resolving step (23) is akin to the idea of resolving an LP in online LP, Bandits with Knapsacks, online allocation problems, where logarithmic or better regret is obtained in a series of work Flajolet and Jaillet (2015), Arlotto and Gurvich (2019), Li and Ye (2022), Li et al. (2021c), Vera and Banerjee (2021), Vera et al. (2021), Bumpensanti and Wang (2020), Jiang et al. (2022). However, since we have identified the optimal basis $\mathcal{I}$ and $\mathcal{J}$ and we stick to this optimal basis to avoid the non-degeneracy assumption, resolving the LP would become resolving the linear equations in (23).
@@ 6.1 High Probability Bound for Identifying One Optimal Basis
We now present the theorem showing that Algorithm 1 indeed helps us identify one optimal basis with a high probability. In practice, the value of $\varepsilon$ will be set to $1/N$ in the following theorem.
@@
For any $\varepsilon>0$, as long as $N_{0}$ satisfies the condition
$2\mathrm{Gap}_{1}(N_{0},\varepsilon)+2\mathrm{Gap}_{2}(N_{0},\varepsilon)\leq\min\{\delta_{1},\delta_{2}\}$ (26)
the outputs $\mathcal{I}_{N_{0}}$ and $\mathcal{J}_{N_{0}}$ of Algorithm 1 satisfy the conditions described in Lemma 3.1 with probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\varepsilon$. Moreover, the sets $\mathcal{I}_{n}$ and $\mathcal{J}_{n}$ will be common for any $n\geq N_{0}$ satisfying (26), which we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$.
The formal proof of Theorem 6.1 is relegated to Section 11. Note that there can be multiple optimal bases to the LP (8), and our approach essentially identifies one particular optimal basis. To see this point, we sort all the elements of the solution $\bm{q}$, and in Algorithm 1, we try to delete each element $i$ from the set $\mathcal{I}$ from the first one to the last one in this order. For a collection of non-zero basic variables denoted as $(i_{1},\dots,i_{M})$, our approaches identifies the collection with $i_{1}$ being the largest among all collections. Then, among all collections with $i_{1}$ being the first element, our approach identifies the collection with $i_{2}$ being the largest, so on and so forth. Our Algorithm 1 identifies such a collection of non-zero basic variables. Similarly, for the redundant constraints, our approach deletes these constraints from the smallest index in the set $[K]$. Therefore, as long as $n$ (which is $2^{m}$ for an integer $m$) is large enough, our Algorithm 1 outputs the same index sets $\mathcal{I}$ and $\mathcal{J}$ that satisfies the conditions in Lemma 3.1, which we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ from now on.
@@
$A^{*}=\left[\begin{aligned} &C(\mathcal{J}^{*},\mathcal{I}^{*})\\ &B(:,\mathcal{I}^{*})\end{aligned}\right].$ (27)
We then denote by $\{\sigma_{1}(A^{*}),\dots,\sigma_{|\mathcal{S}|+K^{\prime}}(A^{*})\}$ the eigenvalues of a matrix $A^{*}$. We define $\sigma$ as
$\sigma=\min\left\{|\sigma_{1}(A^{*})|,\dots,|\sigma_{|\mathcal{S}|+K^{\prime}}(A^{*})|\right\}.$ (28)
From the non-singularity of the matrix $A^{*}$, we know that $\sigma>0$.
Remark. It is clear to see that $\sigma$ in (28) relates to the conditional number of the matrix $A^{*}$. In fact, it holds that $\kappa(A^{*})\leq\frac{|\mathcal{S}|+K}{\sigma}$, where $\kappa(A^{*})$ denotes the conditional number of the matrix $A^{*}$. The value of $\sigma$ will show up in our final bounds as it characterizes how hard it is to learn the matrix $A^{*}$. Also, $\sigma$ gives a natural upper bound of the range of the optimal dual variable corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Note that in previous literature that establishes logarithmic regret for online LP/resource allocation (e.g. Vera and Banerjee (2021), Li et al. (2021c), Jiang et al. (2022)), the conditional number of the constraint matrix will also show up in the final bounds. Therefore, we regard the existence of $\sigma$ as the consequence of adopting the online LP/resource allocation framework. On the other hand, for the MDP problem without resource constraints, the matrix $A^{*}$ can simply be represented as the probability transition matrix and one can easily show that $\sigma\geq 1-\gamma$. We will use this bound when deriving our guarantees for MDP problems without resource constraints.
@@ 6.2 Regret Bounds
We now show that Algorithm 2 achieves a logarithmic regret, as defined in (3). The next lemma shows that to analyze the regret, it suffices to consider the remaining constraint capacities belonging to the binding constraints. The formal proof is relegated to Section 12.
@@
For the optimal basis identified in Algorithm 1 and the corresponding optimal solution $\bm{q}^{*}$, we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ the output sets as long as $N_{0}$ satisfies the condition (26). We also denote by $(\bm{y}^{*},\bm{c}^{*})$ the corresponding optimal dual solution. Then, it holds that
$N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq\sum_{j\in\mathcal{J}^{*}}y^{*}_{j}\cdot\mathbb{E}[\alpha^{N}_{j}]+\sum_{s\in\mathcal{S}}z^{*}_{s}\cdot\mathbb{E}[\mu^{N}_{s}].$ (29)
Therefore, it suffices to analyze how the “remaining resources” $(\bm{\alpha}_{\mathcal{J}^{*}}^{n},\bm{\mu}^{n})$ behave. We now define
$\tilde{\alpha}_{k}(n)=\frac{\alpha^{n}_{k}}{N-n},\leavevmode\nobreak\ \forall k\in\mathcal{J}^{*}\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }\tilde{\mu}_{s}(n)=\frac{\mu^{n}_{s}}{N-n},\leavevmode\nobreak\ \forall s\in\mathcal{S},\forall n\in[N].$ (30)
The key is to show that the stochastic process $\tilde{\alpha}_{k}(n)$ and $\tilde{\mu}_{s}$ possess some concentration properties such that they will stay within a small neighborhood of their initial value $\alpha^{1}_{k}$ and $\mu^{1}_{s}$ for a sufficiently long time. We denote by $\tau$ the time that one of $\tilde{\alpha}_{k}(n)$ for each $k\in\mathcal{J}^{*}$ and $\tilde{\mu}_{s}$ for each $s\in\mathcal{S}$ escape this neighborhood. Then, both $\text{Regret}_{r}(\pi,N)$ and $\text{Regret}_{k}(\pi,N)$ for each $k\in[K]$ can be upper bounded by $\mathbb{E}[N-\tau]$. From the update rule (24) and (25), we know that
$\tilde{\alpha}_{k}(n+1)=\tilde{\alpha}_{k}(n)-\frac{\sum_{(s,a)\in\mathcal{I}^{*}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a)-\tilde{\alpha}_{k}(n)}{N-n-1},\leavevmode\nobreak\ \forall k\in\mathcal{J}^{*}$ (31)
@@
$\tilde{\mu}_{s}(n+1)=\tilde{\mu}_{s}(n)-\frac{\sum_{(s,a)\in\mathcal{I}^{*}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}})-\tilde{\mu}_{s}(n)}{N-n-1}.$ (32)
Ideally, both $\tilde{\alpha}_{k}(n+1)$ and $\tilde{\mu}_{s}(n+1)$ will have the same expectation as $\tilde{\alpha}_{k}(n)$ and $\tilde{\mu}_{s}(n)$ such that they become a martingale. However, this is not true since we have estimation error over $C(\mathcal{J}^{*},\mathcal{I}^{*})$ and $B(:,\mathcal{I}^{*})$, and we only use their estimates to compute $\bm{q}^{n}$. Nevertheless, we can show that $\tilde{\alpha}_{k}(n)$ for each $k\in\mathcal{J}^{*}$ and $\tilde{\mu}_{s}$ for each $s\in\mathcal{S}$ behave as a sub-martingale. Then, from the concentration property of the sub-martingale, we upper bound $\mathbb{E}[\alpha^{N}_{k}]$ for each $k\in\mathcal{J}^{*}$ and $\mathbb{E}[\mu^{N}_{s}]$ for each $s\in\mathcal{S}$. The term $|\mathbb{E}[\alpha^{N}_{k}]|$ for each $k\in[K]\backslash\mathcal{J}^{*}$ can be upper bounded as well. The results are presented in the following theorem. The formal proof is relegated to Section 13 and we provide a brief sketch of the proof for illustration.

### v4

@@ 5.3 Our Formal Algorithm
We now describe our formal algorithm. From the output of Algorithm 2, we characterize one optimal solution. We denote by $\hat{\mathcal{I}}$ and $\hat{\mathcal{J}}$ the output of Algorithm 2. To be specific, we have $\bm{q}^{*}_{\hat{\mathcal{I}}^{c}}=0$ and the non-zero elements $\bm{q}^{*}_{\hat{\mathcal{I}}}$ can be given as the solution to
$\begin{bmatrix}C(\hat{\mathcal{J}}_{1},\hat{\mathcal{I}})\\ B(\hat{\mathcal{J}}_{2},\hat{\mathcal{I}})\end{bmatrix}\cdot\bm{q}^{*}_{\hat{\mathcal{I}}}=\begin{bmatrix}\bm{\alpha}_{\hat{\mathcal{J}}_{1}}\\ \bm{\mu}_{\hat{\mathcal{J}}_{2}}\end{bmatrix}$, (23)
@@
1:Input: the number of rounds $N_{1}$ carrying out Algorithm 2 and the number of rounds $N_{2}$ for resolving.
2:Query the generative model for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ by $N_{1}$ times and obtain the dataset $\mathcal{F}_{N_{1}}$.
3:Obtain the output $\hat{\mathcal{I}}$ and $\hat{\mathcal{J}}$ from Algorithm 2 with input $\mathcal{F}_{N_{1}}$ and $\epsilon=O(\varepsilon^{2}/(K\cdot|\mathcal{S}||\mathcal{A}|))$.
4:Let $\hat{\mathcal{J}}_{1}=\hat{\mathcal{J}}\cap[K]$ and $\hat{\mathcal{J}}_{2}=\hat{\mathcal{J}}\cap\mathcal{S}$.
5:Initialize $\mathcal{H}_{1}=\emptyset$, $\bm{\alpha}^{1}=N_{2}\cdot\bm{\alpha}$ and $\bm{\mu}^{1}=N_{2}\cdot\bm{\mu}$.
6:for $n=1,\dots,N_{2}$ do
@@
8:  Construct a solution $\tilde{\bm{q}}^{n}$ such that $\tilde{\bm{q}}^{n}_{\hat{\mathcal{I}}^{c}}=0$ and $\tilde{\bm{q}}^{*}_{\hat{\mathcal{I}}}$ is the solution to
$\left[\begin{aligned} &\bar{C}^{n}(\hat{\mathcal{J}}_{1},\hat{\mathcal{I}})\\ &\bar{B}^{n}(\hat{\mathcal{J}}_{2},\hat{\mathcal{I}})\end{aligned}\right]\cdot\tilde{\bm{q}}^{n}_{\hat{\mathcal{I}}}=\left[\begin{aligned} &\frac{\bm{\alpha}^{n}_{\hat{\mathcal{J}}_{1}}}{N_{2}-n+1}\\ &\frac{\bm{\mu}^{n}_{\hat{\mathcal{J}}_{2}}}{N_{2}-n+1}\end{aligned}\right].$ (24)
9:  Project $\tilde{\bm{q}}^{n}$ to the set $\{\bm{q}:\|\bm{q}\|_{1}\leq 2,\bm{q}_{\hat{\mathcal{I}}^{c}}=0\}$ to obtain $\bm{q}^{n}$.
10:  For each $(s,a)\in\hat{\mathcal{I}}$, we query the model $\mathcal{M}$ to obtain a sample of the reward $r^{n}(s,a)$ and the costs $c^{n}_{k}(s,a)$ for each $k\in\hat{\mathcal{J}}_{1}\subset[K]$, as well as the state transition $s^{n}(s,a)$.
11:  Update $\mathcal{H}_{n+1}=\mathcal{H}_{n}\cup\left\{r^{n}(s,a),c^{n}_{k}(s,a),s^{n}(s,a),\forall(s,a)\in\hat{\mathcal{I}},\forall k\in\hat{\mathcal{J}}_{1}\right\}$.
12:  Denote by $\bm{c}^{n}(s,a)=(c^{n}_{k}(s,a))_{\forall k\in\hat{\mathcal{J}}_{1}}$ and do the update:
$\bm{\alpha}^{n+1}_{\hat{\mathcal{J}}_{1}}=\bm{\alpha}^{n}_{\hat{\mathcal{J}}_{1}}-\sum_{(s,a)\in\hat{\mathcal{I}}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a).$ (25)
13:  Do the update:
$\bm{\mu}^{n+1}_{s^{\prime}}=\bm{\mu}^{n}_{s^{\prime}}-\sum_{(s,a)\in\hat{\mathcal{I}}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall s^{\prime}\in\hat{\mathcal{J}}_{2}.$ (26)
@@
14:end for
15:Output: $\bm{q}^{1},\dots,\bm{q}^{N_{2}}$.
[…]
Remark 5.3
Note that the adaptive resolving step (24) is akin to the idea of resolving an LP in online LP, Bandits with Knapsacks, online allocation problems, where logarithmic or better regret is obtained in a series of work flajolet2015logarithmic, arlotto2019uniformly, li2022online, li2021symmetry, vera2021bayesian, vera2021online, bumpensanti2020re, jiang2022degeneracy. However, since we have identified the optimal basis $\mathcal{I}$ and $\mathcal{J}$ and we stick to this optimal basis to avoid the non-degeneracy assumption, resolving the LP would become resolving the linear equations in (24).
@@ 6.1 High Probability Bound for Identifying One Optimal Basis
We now present the theorem showing that Algorithm 2 indeed helps us identify one optimal basis with a high probability at least $1-\epsilon$. We specify the value of $\epsilon$ in the following parts.
@@
For any $\epsilon>0$, as long as $N_{0}$ satisfies the condition
$\mathrm{Rad}(N_{0},\epsilon)=\sqrt{\frac{\log(2/\epsilon)}{2N_{0}}}\leq O\left(\min\left\{\delta_{1},\delta_{2},\sigma_{0}\right\}\right)$ (27)
where $\delta_{1}$ defined in (20), $\delta_{2}$ defined in (21) and $\sigma_{0}$ defined in (22), the outputs $\mathcal{I}_{N_{0}}$ and $\mathcal{J}_{N_{0}}$ of Algorithm 2 satisfy the conditions described in Lemma 3.1 with probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\epsilon$. Moreover, denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ the output of Algorithm 1 on the true LP $V$. Then, for any $N_{0}$ satisfying (27), with a probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\epsilon$, we have $\mathcal{I}_{N_{0}}=\mathcal{I}^{*}$ and $\mathcal{J}_{N_{0}}=\mathcal{J}^{*}$.
The formal proof of Theorem 6.1 is relegated to Section 13.1. Note that there can be multiple optimal bases to the LP (8), and our approach essentially identifies one particular optimal basis. To see this point, denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ the output of Algorithm 1 on the true LP $V$. The second part of Theorem 6.1 says that with a high probability, we will have exactly $\mathcal{I}_{N_{0}}=\mathcal{I}^{*}$ and $\mathcal{J}_{N_{0}}=\mathcal{J}^{*}$. Therefore, though the LP (8) can be degenerate and enjoys multiple optimal basis, our Algorithm 2 identifies one particular collection of non-zero basic variables and the supporting constraints. Note that though the bound in Theorem 6.1 depends on the parameters $\delta_{1}$, $\delta_{2}$, and $\sigma_{0}$ that could be unknown, we can actually modify the step 2-3 of our Algorithm 3 a little bit to apply a doubling trick to exponentially increase the sample size that serves as input to Algorithm 2. We can guarantee that the input $N_{1}$ to Algorithm 3 is large enough such that the condition in (27) will be satisfied. Then, Theorem 6.1 guarantees that with a high probability, we find $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ in Algorithm 3, i.e., it holds that $\hat{\mathcal{I}}$$\mathcal{I}^{*}$ and $\hat{\mathcal{J}}=\mathcal{J}^{*}$ with a high probability.
@@
$A^{*}=\left[\begin{aligned} &C(\mathcal{J}_{1}^{*},\mathcal{I}^{*})\\ &B(\mathcal{J}_{2}^{*},\mathcal{I}^{*})\end{aligned}\right]$ (28)
where $\mathcal{J}^{*}_{1}=\mathcal{J}^{*}\cap[K]$ and $\mathcal{J}^{*}_{2}=\mathcal{J}^{*}\cap\mathcal{S}$. We define $\sigma^{*}$ as
$\sigma^{*}=\sigma_{\min}(A^{*})$ (29)
where $\sigma_{\min}(A^{*})$ refers to the smallest singular value of the matrix $A^{*}$. From the non-singularity of the matrix $A^{*}$, we know that $\sigma^{*}>0$.
Remark 6.2
It is clear to see that $\sigma^{*}$ in (29) relates to the conditional number of the matrix $A^{*}$. In fact, it holds that $\kappa(A^{*})\leq\frac{|\mathcal{S}|+K}{\sigma^{*}}$, where $\kappa(A^{*})$ denotes the conditional number of the matrix $A^{*}$. The value of $\sigma^{*}$ will show up in our final bounds as it characterizes how hard it is to learn the matrix $A^{*}$. Also, $\sigma^{*}$ gives a natural upper bound of the range of the optimal dual variable corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Note that in previous literature that establishes logarithmic regret for online LP/resource allocation (e.g. vera2021bayesian, li2021symmetry, jiang2022degeneracy), the conditional number of the constraint matrix will also show up in the final bounds. Therefore, we regard the existence of $\sigma^{*}$ as the consequence of adopting the online LP/resource allocation framework. On the other hand, for the MDP problem without resource constraints, the matrix $A^{*}$ can simply be represented as the probability transition matrix and one can easily show that $\sigma^{*}\geq 1-\gamma$. We will use this bound when deriving our guarantees for MDP problems without resource constraints.
@@ 6.2 Regret Bounds
We now show that Algorithm 3 achieves a logarithmic regret, as defined in (3). The next lemma shows that to analyze the regret, it suffices to consider the remaining constraint capacities belonging to the binding constraints. The formal proof is relegated to Section 13.2.
@@
For the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ identified in Algorithm 2 with a high probability, we denote by $\bm{q}^{*}$ the corresponding primal optimal basic solution. We also denote by $(\bm{y}^{*},\bm{c}^{*})$ the optimal dual solution corresponding to $\bm{q}^{*}$. Then, it holds that
$N_{2}\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N_{2}}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq\sum_{j\in\mathcal{J}_{1}^{*}}y^{*}_{j}\cdot\mathbb{E}[\alpha^{N_{2}}_{j}]+\sum_{s\in\mathcal{J}_{2}^{*}}z^{*}_{s}\cdot\mathbb{E}[\mu^{N_{2}}_{s}],$ (30)
where $\{\bm{q}^{1},\dots,\bm{q}^{N_{2}}\}$ is the output of Algorithm 3.
We now condition on the event that $\hat{\mathcal{I}}$$\mathcal{I}^{*}$ and $\hat{\mathcal{J}}=\mathcal{J}^{*}$. It suffices to analyze how the “remaining resources” $(\bm{\alpha}_{\mathcal{J}_{1}^{*}}^{n},\bm{\mu}_{\mathcal{J}^{*}_{2}}^{n})$ behave. We now define
$\tilde{\alpha}_{k}(n)=\frac{\alpha^{n}_{k}}{N_{2}-n},\penalty 10000\ \forall k\in\mathcal{J}_{1}^{*}\text{\penalty 10000\ and\penalty 10000\ }\tilde{\mu}_{s}(n)=\frac{\mu^{n}_{s}}{N_{2}-n},\penalty 10000\ \forall s\in\mathcal{J}^{*}_{2},\forall n\in[N_{2}].$ (31)
The key is to show that the stochastic process $\tilde{\alpha}_{k}(n)$ and $\tilde{\mu}_{s}$ possess some concentration properties such that they will stay within a small neighborhood of their initial value $\alpha^{1}_{k}$ and $\mu^{1}_{s}$ for a sufficiently long time. We denote by $\tau$ the time that one of $\tilde{\alpha}_{k}(n)$ for each $k\in\mathcal{J}_{1}^{*}$ and $\tilde{\mu}_{s}(n)$ for each $s\in\mathcal{J}^{*}_{2}$ escape this neighborhood. Then, both $\text{Regret}_{r}(\pi,N_{2})$ and $\text{Regret}_{k}(\pi,N_{2})$ for each $k\in[K]$ can be upper bounded by $\mathbb{E}[N_{2}-\tau]$. From the update rule (25) and (26), we know that
$\tilde{\alpha}_{k}(n+1)=\tilde{\alpha}_{k}(n)-\frac{\sum_{(s,a)\in\mathcal{I}^{*}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a)-\tilde{\alpha}_{k}(n)}{N_{2}-n-1},\penalty 10000\ \forall k\in\mathcal{J}_{1}^{*}$ (32)
@@
$\tilde{\mu}_{s}(n+1)=\tilde{\mu}_{s}(n)-\frac{\sum_{(s,a)\in\mathcal{I}^{*}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}})-\tilde{\mu}_{s}(n)}{N_{2}-n-1},\penalty 10000\ \forall s\in\mathcal{J}_{2}^{*}.$ (33)
Ideally, both $\tilde{\alpha}_{k}(n+1)$ and $\tilde{\mu}_{s}(n+1)$ will have the same expectation as $\tilde{\alpha}_{k}(n)$ and $\tilde{\mu}_{s}(n)$ such that they become a martingale. However, this is not true since we have estimation error over $C(\mathcal{J}_{1}^{*},\mathcal{I}^{*})$ and $B(\mathcal{J}^{*}_{2},\mathcal{I}^{*})$, and we only use their estimates to compute $\bm{q}^{n}$. Nevertheless, we can show that $\tilde{\alpha}_{k}(n)$ for each $k\in\mathcal{J}_{1}^{*}$ and $\tilde{\mu}_{s}$ for each $s\in\mathcal{J}_{2}^{*}$ behave as a sub-martingale. Then, from the concentration property of the sub-martingale, we upper bound $\mathbb{E}[\alpha^{N_{2}}_{k}]$ for each $k\in\mathcal{J}_{1}^{*}$ and $\mathbb{E}[\mu^{N_{2}}_{s}]$ for each $s\in\mathcal{J}_{2}^{*}$. The term $|\mathbb{E}[\alpha^{N_{2}}_{k}]|$ for each $k\in[K]\backslash\mathcal{J}_{1}^{*}$, and the term $\mathbb{E}[\mu^{N_{2}}_{s}]$ for each $s\in\mathcal{S}\backslash\mathcal{J}_{2}^{*}$, can be upper bounded as well. The results are presented in the following theorem. The formal proof is relegated to Section 13.3 and we provide a brief sketch of the proof for illustration.